An athlete bicycles an average of 12 mph faster than she runs. She runs 0.8 hours and bicycles for 1.2 hours for a total distance of 30.4 miles. Find her average rate bicycling
Answer by mananth(16949)

An athlete bicycles an average of 12 mph faster than she runs. She runs 0.8 hours and bicycles for 1.2 hours for a total distance of 30.4 miles. Find her average rate bicycling
let her run at x mph
distance run = 0.8x
she rides at x+12 mph
distance she rides = 1.2 ( x+12)
=1.2x + 14.4
0.8x +1.2x+14.4 = 30.4

2x +14.4 = 30.4
2x=30.4-14.4
2x= 16
x= 8
x=8 running speed
riding speed = 20 mph